You and your friends go to a restaurant that has mints in a bowl. As you leave you grab one third of the mints from the bowl. After that your friend Ashely takes 7 mints .Hillary then takes twice as many minutes as Ashley did. There are now nine minutes left in the bowl. How many were there before you and your friend took them
Answer by Boreal(15235)

x mints in bowl
(2/3)x left after you graph a third
(2/3)x-7 after Ashley
(2/3)x-21 after Hillary.
There are 9 left
(2/3)x-21=9
(2/3)x=30
x=45 mints ANSWER
30 were left after you
23 after Ashley
9 after Hillary.